OpenStudy (anonymous):

10^4 √6 + 4^3 √6 - 9^3 √6 -3^4 √6

OpenStudy (anonymous):

10^4=10000

OpenStudy (anonymous):

so 10000x2.4 (sqrt 6=2.4)

OpenStudy (anonymous):

10000x2.4=24000

OpenStudy (anonymous):

ok i did that part wrong

OpenStudy (anonymous):

64x2.4=153.6

OpenStudy (anonymous):

24000+153.6=24153.6

OpenStudy (anonymous):

now we work on the other side

OpenStudy (anonymous):

9^3=729 729x2.4=1749.6

OpenStudy (anonymous):

now one thing i dont understand; is this subraction or is this a negative number? "-3^4"

OpenStudy (anonymous):

i just want to be sure because if its negative its 81

OpenStudy (anonymous):

(Assuming its negative) 81x2.4=194.4 194.4x1794.6=348870.24 so 348870.24-24153.6=324716.64